OpenStudy (anonymous):

There were 150 freshman and sophomores at the high school musical Reese and her friends attended. Freshman tickets were $2.00 each and sophomore tickets were $3.00 each. The receipts for freshman and sophomore tickets totaled $390. How many sophomores bought tickets? 100 90 60 30

OpenStudy (anonymous):

@shaik0124

OpenStudy (anonymous):

@shaik0124

OpenStudy (king):

let the no. of freshmen be x, and no. of sophomores be y, then x+y=150 x=150-y also, 2x+3y=390 2(150-y)+3y=390 300-2y+3y=390 y=90

OpenStudy (king):

i.e no of sophomores is 90....:D